Obtain the Conjunctive Normal Form of (x^y) V (-x^y)




1
Expert's answer
2022-01-10T07:44:59-0500

"xy+x'y=(xy+x')(xy+y)=(x+x')(x'+y)(x+y)(y+y)="


"=(x'+y)(x+y)y"